A body is projected vertically upwards with velocity u. It crosses a point
in its journey at a height h twice just after 1 and 7 seconds. The value
of u in m/s is (g=10m/s 2 )

Grade:11

1 Answers

Lakshmipathi
7 years ago
HAI PRAVEEN
 
As the body is projected upwards we can write the equation as
h=ut-1/2gt2
gt2 -2ut +2h=0
t1+t2 are roots of equation
therefore t1+t2=[2u/g] ----->u=g/2(t1+t2)
therfore u=40 ms-1 or 40 m/s both are same.